OpenStudy (anonymous):

What is the axis of symmetry for y = {x^2} - 8x + 8? Be sure to use correct notation.

OpenStudy (alekos):

axis of symmetry, x = -b/2a

OpenStudy (alekos):

where the general form of a parabola is y= ax^2+bx+c

OpenStudy (anonymous):

Okay I know the vertex is 4,-8

OpenStudy (alekos):

well then the axis of symmetry is just the x value of the vertex

OpenStudy (anonymous):

which is 4 my other problem was setting it up as an equation

OpenStudy (anonymous):

y = (x-4)^2 - 8 this is what i have

OpenStudy (alekos):

yeah thats correct

OpenStudy (anonymous):

though it says I was wrong when I tried to submit it

OpenStudy (alekos):

who says it was wrong?

OpenStudy (anonymous):

when I try to submit my assignment it counts the equation wrong

OpenStudy (alekos):

But that is the correct vertex form of your original equation and x=4 is the equation for the axis of symmetry

OpenStudy (anonymous):

that's what happen I have -4

OpenStudy (alekos):

ok, there you go

OpenStudy (anonymous):

so its x=4

OpenStudy (alekos):

oh yes it is

OpenStudy (anonymous):

it's right thank you :D

OpenStudy (alekos):

no problemo

OpenStudy (anonymous):

although I have more questions I may need help on in a few minutes